. Calculate the price elasticity of the following products. State whether elasticity of supply is elastic, unit elastic, or inelastic Cocoa Puffs: The price of a 14-oz. box of Cocoa Puffs rises 4 percent and the quantity supplied rises 15 percent. a. b. Japanese yen: The price of Japanese yen in terms of dollars rises from 100 yen per dollar to 110 yen per dollar. Its quantity supplied rises from 5,000,000 yen to 5,300,000 yen per year Jansport backpacks: The price of Jansport backpacks falls from $30 a pack to $25 a pack. The quantity supplied falls from 150,000 to 125,000 per week. c.
Answer #1

Price elasticity is calculated using the following formula:

Price elasticity = % change in quantity /% change in price.

A. Price elasticity = % change in quantity/% change in price = 15/4 = 3.75

Also, the elasticity of supply is 'elastic' because the value of elasticity of supply is greater than 1.

B. % change in price = (110-100)/100 *100 = 10/100 *100 = 10%

% change in quantity = (5,300,000 - 5,000,000)/5,000,000 * 100 = 300,000/5,000,000 * 100 = 3/50*100 = 6%

Price elasticity = % change in quantity/% change in price = 6/10 = 0.6

Also, the elasticity of supply is 'inelastic' because the value of elasticity of supply is less than 1.

C.

% change in price = (25-30)/25 *100 = -5/25 *100 = -20%

% change in quantity = (125,000 - 150,000)/125,000 * 100 = -25,000/150,000 * 100 = -1/6*100 = -16.67%

Price elasticity = % change in quantity/% change in price =-16.67/-20 = 0.83

Also, the elasticity of supply is 'inelastic' because the value of elasticity of supply is less than 1.